OpenStudy (anonymous):

A city has a population of 320,000 people. Suppose that each year the population grows by 7.25%. What will the population be after 12 years?

OpenStudy (gorv):

you know the formula for calculating this ???

OpenStudy (anonymous):

I=prt

OpenStudy (gorv):

this is for simple .......for compound???

OpenStudy (gorv):

population after 12 year = \[P*(1+\frac{ r}{ 100 })^{t}\]

OpenStudy (gorv):

use this

OpenStudy (anonymous):

r will be what ?

OpenStudy (anonymous):

i got 2.67

OpenStudy (gorv):

r=7.25 p=320000 t=12

OpenStudy (anonymous):

is my answer correct

OpenStudy (gorv):

no

OpenStudy (gorv):

\[320000*(1+\frac{ 7.25 }{ 100 })^{12}\]

OpenStudy (gorv):

12 is the power of bracket

OpenStudy (anonymous):

343200^12 is 2.67

OpenStudy (anonymous):

so iit will be 3

OpenStudy (gorv):

\[32000*(\frac{ 100+7.15 }{ 100})^{12}\]

OpenStudy (gorv):

\[32000*(\frac{ 107.25 }{ 100 })^{12}\]

OpenStudy (anonymous):

where does the extra 100 come from .... can i just have the ANSWER PLEASE

OpenStudy (gorv):

\[320000*(1.0725)^{12}\]

OpenStudy (gorv):

no direct ans buddy sorry .....try

OpenStudy (anonymous):

741169.58

OpenStudy (anonymous):

wich stills sounds incorrect

OpenStudy (gorv):

\[1+\frac{ 7.25 }{ 100}=\frac{ 100+7.25 }{ 100 }\]

OpenStudy (gorv):

that's right

OpenStudy (anonymous):

round to the nearest whole number

OpenStudy (gorv):

741169 or 741170

OpenStudy (anonymous):

OMG ! Thought this will never be over !!

OpenStudy (anonymous):

thank u

OpenStudy (gorv):

wlcome :)